# Data Poisoning for In-context Learning

### v2

@@ Abstract
In the domain of large language models (LLMs), in-context learning (ICL) has been recognized for its innovative ability to adapt to new tasks, relying on examples rather than retraining or fine-tuning. This paper delves into the critical issue of ICL’s susceptibility to data poisoning attacks, an area not yet fully explored. We wonder whether ICL is vulnerable, with adversaries capable of manipulating example data to degrade model performance. To address this, we introduce ICLPoison, a specialized attacking framework conceived to exploit the learning mechanisms of ICL. Our approach uniquely employs discrete text perturbations to strategically influence the hidden states of LLMs during the ICL process. We outline three representative strategies to implement attacks under our framework, each rigorously evaluated across a variety of models and tasks. Our comprehensive tests, including trials on the sophisticated GPT-4 model, demonstrate that ICL’s performance is significantly compromised under our framework. These revelations indicate an urgent need for enhanced defense mechanisms to safeguard the integrity and reliability of LLMs in applications relying on in-context learning.
Machine Learning, ICML
@@ 1 Introduction
In recent years, In-Context Learning (ICL) (Brown et al., 2020; Min et al., 2022) has emerged as an important component of large language models (LLMs). Unlike traditional machine learning algorithms that require extensive retraining or fine-tuning to adapt new data (Hoi et al., 2021; Zhang & Yang, 2021; Zhuang et al., 2020), ICL enables LLMs to make predictions based on extra information in the prompt (e.g. a few examples related to a specific task), without changing the model parameters. For example, consider the task of predicting a person’s nationality and the prompt consists of examples, e.g. “Albert Einstein was German; Isaac Newton was English;”, followed by the query “Thomas Edison was”, an LLM such as GPT-4 will predict “American” accurately. The efficiency and flexibility of ICL have gained significant attention and revolutionized various real-world applications, ranging from reasoning (e.g. chain-of-thought (Wei et al., 2022; Wang et al., 2022)) to retrieval-augmented-generation (RAG) (Lewis et al., 2020).
The success of ICL depends critically on the examples it utilizes. Studies have shown that the ICL performance is sensitive to certain characteristics of demonstrations, e.g., the selection of examples (Wang et al., 2023) and the order of examples in demonstration (Min et al., 2022). Therefore, it naturally raises a question: Is ICL vulnerable to potential data poisoning attacks? In reality, malicious actors may manipulate demonstration data to degrade the model performance. For example, adversaries can execute data poisoning attacks by strategically altering the examples in demonstrations used in ICL. Such attacks aim to disrupt the learning process, leading to inaccurate or biased predictions.
@@
In this paper, we aim to answer the above question by delving into data poisoning attacks in ICL and uncovering the vulnerability of ICL against these attacks. We consider the standard pipeline of ICL, where examples are randomly selected from a certain data set to tailor ICL for a downstream task. In terms of the attack, we assume that the adversary deliberately alters some examples in the data, and his goal is to ensure that the learning process is adversely affected and the overall performance decreases. This scenario is both significant and practical. For instance, in a RAG system where demonstrations are retrieved from a dataset, the attacker manipulates some data from a specific domain or a task (e.g. reviews about a specific brand) to downgrade the quality of the generated content of LLM related to that domain/task. To the best of our knowledge, we are the first to investigate the vulnerability of ICL against data poisoning attacks.
@@
Data poisoning in ICL faces both unique challenges specific to ICL and common obstacles in traditional data poisoning. First, in contrast to traditional learning algorithms with explicit training objectives, ICL enables LLMs to learn from demonstrations without explicit optimization (Brown et al., 2020; Min et al., 2022). Since traditional poisoning strategies are designed specifically to target the training process and exploit loss functions in conventional models (Biggio et al., 2012; Steinhardt et al., 2017; Geiping et al., 2020; He et al., 2023), the discrepancy between ICL and these strategies implies that they are not directly applicable to ICL. Conducting effective data poisoning attacks for ICL demands a thorough understanding of the unique learning mechanism of ICL. Second, similar to traditional attacking methods, data poisoning for ICL also requires creating samples that are imperceptible to humans yet disruptive. These poisoned examples must seamlessly integrate with the other data to harm the learning process in a subtle way. While both traditional data poisoning and the attack in ICL require careful designs of the attack, for ICL, one extra challenge arises from the discrete vocabulary of language models, making it hard to manipulate inputs for effective disturbance (Lei et al., 2019; Xu et al., 2023).
@@
To tackle the above challenges, we introduce a novel and versatile attacking framework, ICLPoison, to exploit the unique learning mechanism of ICL. In particular, previous research (Xie et al., 2021; Hendel et al., 2023; Liu et al., 2023b; Wang et al., 2023) has shown a strong correlation between ICL performance and the hidden states within LLMs. Our framework, ICLPoison, ingeniously distorts these hidden states through strategic text perturbations—a non-trivial adaption given the attacker’s limited ability to only alter the examples in the demonstration. We further design three strategies for instantiating and optimizing poisoning attacks under the ICLPoison framework. Comprehensive experiments across various LLMs and tasks demonstrate the effectiveness of our methods, highlighting the vulnerability of ICL. Notably, we have successfully degraded the performance of ICL in advanced models, including GPT-4 (a 10% decrease in ICL accuracy). Our study significantly advances the understanding of ICL’s vulnerabilities to data poisoning, bolstering the security and reliability of LLMs.
@@ 2 Preliminary
In this section, we introduce some background of ICL. We then present the preliminary results that the distortion in hidden states indeed affects the ICL performance, which inspires our framework. Key notations are also introduced.
2.1 In-context Learning (ICL)
[…]
2.2 Understanding of ICL via Hidden States
The emergence of ICL ability in LLMs has gained substantial attention, particularly in understanding its working mechanisms (Xie et al., 2021; Hendel et al., 2023; Von Oswald et al., 2023; Garg et al., 2022; Bai et al., 2023). Researchers have demonstrated that during ICL, LLMs can effectively extract “latent concepts” (denoted as $\theta$) from demonstrations (Xie et al., 2021; Wang et al., 2023), and the ICL predicting procedure can be decomposed as
$P(y^{query}|[S,x^{query}])$ (1)
[…]
where $P(\cdot)$ denotes the probability distribution during the generation process of LLMs. In other words, ICL learns $\theta$ from the demonstration first and then make a prediction for $x^{query}$. The latent concept $\theta$ characterizes the intricate patterns and contextual information embedded in the data regarding the task. A key in this process is the hidden states, represented as $h$, which are defined as the representations of the last token of the input prompt at different layers of the model (as indicated by various studies (Hendel et al., 2023; Liu et al., 2023b; Todd et al., 2023)). The hidden states are used in encoding the latent concepts. In particular, consider a model $f$ composed of $L$ transformer layers, where each layer produces a vector representation $h_{l}(p,f)\in\mathbb{R}^{d}$ for the last token of the input prompt $p$ and $l\in[L]$. It has been observed that the hidden state $h_{l}$ at certain intermediate layers effectively works as $\theta$ (Hendel et al., 2023). Moreover, further research (Liu et al., 2023b; Todd et al., 2023) has been conducted to distill more concise representations from the set of hidden states $\{h_{l}\}_{l=1}^{L}$, enhancing the concept formation in ICL.
2.3 Preliminary Experiments
As discussed in the above, hidden states of the input prompt encode the latent concept and are closely related to the performance of ICL. Therefore, we conduct preliminary experiments to explore whether directly perturbing these hidden states can degrade ICL, and whether ICL is more vulnerable when perturbing the hidden states in all layers.
[…]
Figure 1: Results when directly perturbing the hidden states with Gaussian noises of various standard deviations (noise strength).
3 A Novel Framework: ICLPoison
In this section, we introduce a novel framework, ICLPoison, to conduct data poisoning by distorting hidden states of LLMs during the ICL process.
@@ 3.1 Threat Model
We assume that attackers have the full access to a portion of the dataset, specifically a subset related to a particular task. However, they may lack complete knowledge of the entire prompting set. Crucially, they are unaware of the details of the ICL process, including the test data, the Large Language Models (LLMs) employed, and specific ICL configurations like the number of examples and the templates used for demonstrations. Despite these limitations, the attackers can leverage open-source LLMs to introduce poisoned data into the subset they control.
@@ 3.2 ICLPoison Design
As highlighted in Section 2.1, ICL differs from traditional learning algorithms since it lacks an explicit training objective that can be directly targeted by data poisoning attacks. To address this unique challenge, we draw inspiration from the dynamics of hidden states from Section 2.2, which indicates that the representation of the last token of the demonstration encodes the latent concept about the task. Building on this understanding, we introduce ICLPoison, a novel data poisoning framework specifically designed for the ICL process. ICLPoison strategically alters examples in demonstrations to distort the hidden states in the LLM, leveraging these internal dynamics to achieve our poisoning goals.
@@
The details of ICLPoison are as follows. We focus on a surrogate LLM $f$ with $L$ layers and a task-specific prompt set $D^{p}_{t}=\{(x_{i,t}^{p},y_{i,t}^{p})\}_{i=1}^{N}$ from task $t$. Our objective is to diminish the ICL accuracy for the task $t$ by maximizing the distortion of hidden states in samples from $D^{p}_{t}$. We propose perturbing the input $x^{p}_{i,t}\sim\mathcal{X}_{t}$ using a transformation $\delta:\mathcal{X}_{t}\rightarrow\mathcal{X}_{t}$ while keeping the label $y^{p}_{i,t}$ unchanged. The perturbation $\delta$ must be imperceptible to humans, hence it is constrained within a set $\Delta$ of imperceptible mappings. Details of $\delta$ will be discussed in Section 3.3. For each example $(x_{i,t}^{p},y_{i,t}^{p})$ in $D^{p}_{t}$, we extract its hidden states. Since the attacker lacks the knowledge of the test data, we use a dummy query $x^{query}_{t}\sim\mathcal{X}_{t}$ as a stand-in (as mentioned in (Hendel et al., 2023)). We then concatenate $(x_{i,t}^{p},y_{i,t}^{p})$ with $x^{query}_{t}$ to create a demonstration, and denote $h_{l}(x_{i,t}^{p},f)$ as the representation of the last token in the $l^{th}$ layer of model $f$. Since our focus is on perturbing $x^{p}_{i,t}$, we omit $y^{p}_{i,t}$ in the following discussion. The representations from all $L$ layers of model $f$ regarding input $x_{i,t}^{p}$ are denoted as $H(x_{i,t}^{p},f):=\{h_{l}(x_{i,t}^{p},f)\}_{l=1}^{L}$, representing the hidden states for $x^{p}_{i,t}$ under model $f$. For the perturbed input, the hidden states are $H(\delta_{i}(x_{i,t}^{p}),f)$, with $\delta_{i}$ being the specific perturbation for $x_{i,t}^{p}$.
@@
Our observations in Figure 1 indicate that altering the hidden state of just one layer is not optimal to significantly impact ICL performance. Therefore, we aim to maximize the minimum difference across all layers between the original and perturbed hidden states. To normalize differences across layers with varying scales, we use the normalized $L_{2}$ norm to measure the distance of the hidden state between the original example and the perturbed one for each layer: $l_{d}(h_{l}(x_{i,t}^{p},f),h_{l}(\delta_{i}(x_{i,t}^{p}),f))=\|\frac{h_{l}(x_{i,t}^{p},f)}{\|h_{l}(x_{i,t}^{p},f)\|_{2}}-\frac{h_{l}(\delta_{i}(x_{i,t}^{p}),f)}{\|h_{l}(\delta_{i}(x_{i,t}^{p}),f)\|_{2}}\|_{2}$. The distortion between $x_{i,t}^{p}$ and $\delta_{i}(x_{i,t}^{p})$ is further defined as:
$\mathcal{L}_{d}(H(x_{i,t}^{p},f),H(\delta_{i}(x_{i,t}^{p}),f))$ (2)
$=\min\limits_{l\in[L]}l_{d}(h_{l}(x_{i,t}^{p},f),h_{l}(\delta_{i}(x_{i,t}^{p}),f)).$
The attacking objective of ICLPoison then becomes
$\max\limits_{\delta_{i}\in\Delta}\mathcal{L}_{d}(H(x_{i,t}^{p},f),H(\delta_{i}(x_{i,t}^{p}),f)).$ (3)
@@
In other words, $\mathcal{L}_{d}(H(x_{i,t}^{p},f),H(\delta_{i}(x_{i,t}^{p}),f))$ denotes the minimum changes (or lower bound of the distortion) caused by the perturbation $\delta_{i}$ across all the layers of the model. This approach ensures that the perturbation $\delta_{i}$ introduces the most substantial change to the hidden states in the LLM during the ICL process, potentially enhancing the effectiveness of the poisoning attack.

### v3

@@ Abstract
In-context learning (ICL) has emerged as a capability of large language models (LLMs), enabling them to adapt to new tasks using provided examples. While ICL has demonstrated its strong effectiveness, there is limited understanding of its vulnerability against potential threats. This paper examines ICL’s vulnerability to data poisoning attacks. We introduce ICLPoison, an attacking method specially designed to exploit ICL’s unique learning mechanisms by identifying discrete text perturbations that influence LLM hidden states. We propose three representative attack strategies, evaluated across various models and tasks. Our experiments, including those on GPT-4, show that ICL performance can be significantly compromised by these attacks, highlighting the urgent need for improved defense mechanisms to protect LLMs’ integrity and reliability.
Data Poisoning for In-context Learning
Pengfei He1thanks: Corresponding to hepengf1@msu.edu, Han Xu2, Yue Xing1, Hui Liu1, Makoto Yamada3, Jiliang Tang1 1Michigan State University  2University of Arizona 3 Okinawa Institute of Science and Technology
@@ 1 Introduction
In-context learning (ICL) (Brown et al., 2020; Min et al., 2022) has emerged as an important capability of large language models (LLMs). Unlike traditional machine learning algorithms that require extensive retraining or fine-tuning to adapt to new tasks (Hoi et al., 2021; Zhang and Yang, 2021; Zhuang et al., 2020), ICL enables LLMs to make predictions based on a few examples related to a specific task, without changing the model parameters. For example, consider the task of predicting a person’s nationality and the prompt consists of examples, e.g. “Albert Einstein was German; Isaac Newton was English;”, followed by the query “Thomas Edison was”, an LLM such as GPT-4 will predict “American” accurately. The efficiency and flexibility of ICL have drawn significant attention and revolutionized various real-world applications, ranging from LLM-based few-shot healthcare agents (Shi et al., 2024) to knowledge tagging in the education domain (Li et al., 2024), where demonstrations are utilized to handle specific tasks.
Despite the success of ICL, studies have shown that the ICL performance is sensitive to certain characteristics of demonstrations, such as the example selection strategy and the quality of examples (Wang et al., 2023; Min et al., 2022). Therefore, it naturally raises a question: Is ICL vulnerable to potential data poisoning attacks? Unlike the traditional data poisoning attack (Steinhardt et al., 2017) where attackers manipulate the training data to corrupt the model trained on the poisoned data, we focus on the perturbations in examples utilized in ICL while keeping the LLMs untouched.
@@
This paper aims to answer the above question by exploring data poisoning attacks in ICL and uncovering the vulnerability of ICL when facing these attacks. We consider the standard pipeline of ICL, where examples are randomly selected from a data set for the corresponding downstream task. In terms of the attack, we assume that the adversary deliberately inserts some poisoned examples in this data, and their goal is to ensure that the learning process is adversely affected and the overall ICL prediction performance decreases. This scenario can be both significant and practical. For instance, LLMs and ICL are used in medical systems (Shi et al., 2024) for multiple tasks including diagnosis and cost computation. Healthcare providers such as hospitals and physicians may manipulate electronic health records (EHR) to increase revenue (Finlayson et al., 2018).
@@
Technically, data poisoning in ICL faces both unique challenges specific to ICL and common obstacles in traditional data poisoning. First, in contrast to traditional learning algorithms with explicit training objectives, ICL enables LLMs to learn from demonstrations without explicit training (Brown et al., 2020; Min et al., 2022). Since traditional poisoning strategies are designed specifically to target the training process and exploit loss functions in conventional models (Biggio et al., 2012; Steinhardt et al., 2017; Geiping et al., 2020; He et al., 2023), they are not directly applicable to ICL. Conducting effective data poisoning attacks for ICL requires a thorough understanding of the unique learning mechanism of ICL. Second, similar to traditional attacking methods, data poisoning for ICL also requires creating samples that are imperceptible to humans yet disruptive. These poisoned examples must seamlessly integrate with the other data to harm the learning process subtly. Moreover, one extra challenge of poisoning ICL arises from the discrete vocabulary of language models, making it hard to manipulate inputs for effective disturbance (Lei et al., 2019; Xu et al., 2023).
@@
To tackle the above challenges, we introduce a novel and versatile attacking method, ICLPoison, to exploit the unique learning mechanism of ICL. In particular, previous research (Xie et al., 2021; Hendel et al., 2023; Liu et al., 2023; Wang et al., 2023) has shown a strong correlation between ICL performance and the hidden states within LLMs. ICLPoison distorts these hidden states through strategic text perturbations, despite the attacker’s limited ability to alter only the demonstration examples. We further design three strategies for instantiating and optimizing poisoning attacks under the ICLPoison method. Extensive and comprehensive experiments across various LLMs and tasks demonstrate the effectiveness of our methods, highlighting the vulnerability of ICL. Notably, we have successfully degraded the performance of ICL in advanced models, including GPT-4 (a 10% decrease in ICL accuracy). Our study enhances the understanding of ICL’s vulnerability to data poisoning, which helps improve the security and reliability of LLMs.
2 Related Works
2.1 Data Poisoning Attack
Data poisoning attacks (Biggio et al., 2012; Steinhardt et al., 2017) usually occur during the data collection phase of machine learning model training, where the training data is tampered with to induce malicious behaviors in the resulting models (He et al., 2023). These malicious behaviors include causing a poisoned model to have a poor overall accuracy (Steinhardt et al., 2017; Fowl et al., 2021; Huang et al., 2021), misclassifying a specified subset of test samples (Shafahi et al., 2018; Zhu et al., 2019), or inserting backdoors (Chen et al., 2017; Gu et al., 2019). The data poisoning on traditional NLP models is widely explored. CARA (Chan et al., 2020) inserts backdoors into text classifiers by generating poisoned samples by utilizing a conditional adversarially regularized autoencoder; (Yang et al., 2021; Schuster et al., 2020) poison the embeddings space of NLP models; (Chen et al., 2021) investigates different triggers to evaluate the effectiveness of backdoor attacks on NLP tasks.
With the development of LLMs, poisoning attacks have also drawn the attention of the whole community, and studies have examined the vulnerability of LLMs against corruption in the pre-training and fine-tuning data. For example, BadPrompt (Cai et al., 2022) inserts backdoors during prompt-tuning and selects effective triggers to maximize the poisoning effect; POISONPROMPT (Yao et al., 2024) focuses on a similar setting and leverages a bi-level optimization objective to optimize the trigger while maintaining utility on clean samples; NOTABLE (Mei et al., 2023) considers the transferability of backdoors and injects backdoors into the encoder to adapt the attack to different downstream tasks and prompting strategies. (Wan et al., 2023) further explores the threat during instruction tuning of LLMs and shows that attackers can manipulate model predictions whenever a desired trigger phrase appears in the input. However, these attacks still focus on altering the pre-training or fine-tuning data, and the threat of poisoning examples in ICL remains unexplored.
2.2 Attacking In-context Learning (ICL)
ICL is considered an efficient way to adapt downstream tasks with a few examples, thus the potential safety issue with regard to ICL is of great importance. For instance, (Qiang, 2024) optimizes adversarial suffixes of examples to mislead the generation of LLMs; (Wei et al., 2023) also employs harmful demonstrations to induce LLMs to produce harmful responses. (Kandpal et al., 2023) first explores the threat of backdoor attacks on ICL. In particular, they insert backdoors in pre-training LLMs via fine-tuning on a poisoned dataset. Then when the prompt consists of examples from the target task, the backdoored LLM will misclassify the test sample containing the trigger as the malicious label. Following works (Zhao et al., 2024a, 2023) propose to directly add triggers into the demonstrations to avoid the computation cost of pre-training or fine-tuning. However, these works require the insertion of obvious triggers that can be easy to detect. In this work, we study more practical scenarios in which imperceptible perturbations are added to demonstrations, and attackers aim to compromise the ICL’s overall effectiveness. This presents unique challenges due to the implicit learning mechanism of ICL.
@@ 3 Preliminary
In this section, we introduce ICL and its hidden states, along with the necessary notations.
ICL. ICL is a paradigm that allows LLMs to learn tasks given a few examples in the form of demonstrations (Brown et al., 2020). To mathematically define the notation of ICL, suppose that we have a pre-trained LLM $f$, which takes in an input prompt $p$ and outputs a response $y$, i.e. $f(p)=y$ [^1] Since the main focus of this paper is not on the generation of LLMs, we adopt the default generation scheme for each LLM.. Given a task $t\in\mathcal{T}$ from the ICL task set $\mathcal{T}$, we assume $(x,y)\sim\mathcal{D}_{t}$ where $x\in\mathcal{X}_{t}$ and $y\in\mathcal{Y}_{t}$. We further assume a prompt set $D_{t}=\{(x_{i,t},y_{i,t})\}_{i=1}^{N}$, where $(x_{i,t},y_{i,t})\sim\mathcal{D}_{t}$. To conduct an ICL prediction for a query $x^{query}\in\mathcal{X}_{t}$ under task $t$, the user first randomly selects $k$ input-output pairs from $D_{t}$ and concatenates them as a demonstration $S$, i.e. $S=[(x_{i,t},y_{i,t})]_{i=1}^{k}$. The demonstration is combined with query $x^{query}$ as an input prompt, and this prompt is sent to the LLM. We define the prediction of ICL as $\hat{y}^{query}_{ICL}=f([S,x^{query}])$. In this work, we consider classification tasks such as sentiment analysis and text classification.
Hidden states of ICL. Extensive studies are conducted to understand the mechanisms of ICL (Xie et al., 2021; Hendel et al., 2023; Von Oswald et al., 2023; Garg et al., 2022; Bai et al., 2023). Researchers have demonstrated that the hidden states (represented as $h$), which are defined as the representations of the last token of the input prompt at different layers of the model (as indicated by various studies (Hendel et al., 2023; Liu et al., 2023; Todd et al., 2023)), plays a critical role in ICL. The hidden states can encode the latent concepts in the examples corresponding to the task, which further guides the prediction. In particular, consider a model $f$ composed of $L$ transformer layers, where each layer produces a vector representation $h_{l}(p,f)\in\mathbb{R}^{d}$ for the last token of the input prompt $p$ and $l\in[L]$. It has been observed that the LLMs can make correct predictions conditioning on hidden states $h_{l}$ extracted from certain intermediate layers (Hendel et al., 2023). The hidden states of LLMs provide numerical representations and condense the information from the examples (Liu et al., 2023; Todd et al., 2023), which inspires our design of ICLPoison, as illustrated in the next section.
4 ICLPoison
In this section, we introduce a novel attacking method, ICLPoison, to conduct data poisoning by distorting hidden states of LLMs during ICL.
@@ 4.1 Threat Model
We assume the attacker’s goal is to compromise the ICL performance when adapting the downstream tasks using examples from the poisoned dataset as the demonstrations. We assume the attacker can insert poisoned data into the dataset. Crucially, attackers are unaware of the details of the ICL process, including the test data, the Large Language Models (LLMs) employed, and specific ICL configurations such as the number of examples and the templates used for demonstrations. Despite these limitations, the attackers can leverage open-source LLMs to generate poisoned data. In real practice, there are many possible scenarios where the attack can insert the poisoned data into the target data set. For instance, attackers can insert misinformation into the database used in a system where data is collected from various sources (Zou et al., 2024); the attackers can also manipulate third-party APIs to insert poisoned examples into the demonstrations (Zhao et al., 2024b). These threats bring safety and ethical concerns for the real applications of ICL in safety-critic domains like healthcare (Shi et al., 2024; Joe et al., 2021), finance (Loukas et al., 2023; Paladini et al., 2023), education (Li et al., 2024).
@@ 4.2 ICLPoison Design
As highlighted in Section 3, ICL differs from traditional learning algorithms since it lacks an explicit training objective that can be directly targeted by data poisoning attacks. To address this unique challenge, we draw inspiration from the understanding of hidden states in Section 3 and introduce ICLPoison, a novel data poisoning attack specifically designed for the ICL process. ICLPoison strategically alters examples in demonstrations to distort the hidden states for the poisoning goal.
@@
The details of ICLPoison are as follows. We focus on a surrogate LLM $f$ with $L$ layers and $(x,y)\sim\mathcal{D}_{t}$. Our objective is to reduce the ICL accuracy for the task $t$ by maximizing the distortion of hidden states of $(x,y)$. We propose perturbing the input $x\sim\mathcal{X}_{t}$ using a transformation $\delta:\mathcal{X}_{t}\rightarrow\mathcal{X}_{t}$ while keeping the label $y$ unchanged. The perturbation $\delta$ must be imperceptible to humans, hence it is constrained within a set $\Delta$ of imperceptible mappings. Details of $\delta$ will be discussed in Section 4.3. For each example $(x,y)$ possessed by the attacker, we extract its hidden states. Since the attacker lacks the knowledge of the test data, we use a dummy query $x^{query}_{t}\sim\mathcal{X}_{t}$ as a stand-in as suggested in (Hendel et al., 2023). We then concatenate $(x,y)$ with $x^{query}_{t}$ to create a demonstration and denote $h_{l}(x,f)$ as the representation of the last token in the $l^{th}$ layer of model $f$. Since our focus is on perturbing $x$, we omit $y$ in the following discussion. The representations from all $L$ layers of model $f$ regarding input $x$ are denoted as $H(x,f):=\{h_{l}(x,f)\}_{l=1}^{L}$, representing the hidden states for $x$ under model $f$. For the perturbed input, the hidden states are $H(\delta_{x}(x),f)$, with $\delta_{x}$ being the specific perturbation for $x$.
@@
To maximize the poisoning effect, we aim to maximize the minimum difference across all layers between the original and perturbed hidden states. To normalize differences across layers with varying scales, we use the normalized $L_{2}$ norm to measure the distance of the hidden state between the original example and the perturbed one for each layer: $l_{d}(h_{l}(x,f),h_{l}(\delta_{x}(x),f))=\|\frac{h_{l}(x,f)}{\|h_{l}(x,f)\|_{2}}-\frac{h_{l}(\delta_{x}(x),f)}{\|h_{l}(\delta_{x}(x),f)\|_{2}}\|_{2}$. The distortion between $x$ and $\delta_{i}(x)$ is further defined as:
$\mathcal{L}_{d}(H(x,f),H(\delta_{x}(x),f))$ (1)
$=\min\limits_{l\in[L]}l_{d}(h_{l}(x,f),h_{l}(\delta_{x}(x),f)).$
The attacking objective of ICLPoison is
$\max\limits_{\delta_{x}\in\Delta}\mathcal{L}_{d}(H(x,f),H(\delta_{x}(x),f)).$ (2)
@@
In other words, $\mathcal{L}_{d}(H(x,f),H(\delta_{x}(x),f))$ denotes the minimum changes (or lower bound of the distortion) caused by the perturbation $\delta_{x}$ across all the layers of the model. This approach ensures that the perturbation $\delta_{x}$ introduces the most substantial change to the hidden states in the LLM during ICL. By optimizing the objective in Equation 2 for each accessible example $(x,y)$, the attacker can create a poisoned example set $D^{p}_{t}$.